Circular No. 6748

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

S5 0716+71 AND PKS 0422+004 G. Tosti, Cattedra di Astrofisica, Università di Perugia, reports: “Ob-

servations with the 0.4-m Robotic Telescope at the Perugia University Observatory (Tosti et al. 1996, PASP 108, 706) show both S5 0716+71 (Kuhr et al. 1981, A.J. 86, 854) and PKS 0422+004 to be in a major outburst with Cousins R Cousins magnitudes as follows: S5 0716+71, Sept. 9 UT, 12.99 ± 0.05; 10, 13.12 ± 0.04; 18, 12.69 ± 0.03; 19, 12.58 ± 0.03. PKS 0422+004, Sept. 10, 13.46 ± 0.04; 18, 13.34 ± 0.04; 19, 13.42 ± 0.04; 20, 13.30 ± 0.04. The brightness of S5 0716+71 increased by more than 1 mag with respect to the mean value observed during the last three years (R = 13.9). The magnitudes of PKS 0422+004 are very close to those observed at the peak of the Nov. 1994 historic maximum (R = 13.2; V = 13.65; e.g., IAUC 6106; Massaro et al. 1996, A.Ap. 314, 87). Further observations are encouraged at all wavelengths.”

XTE J1739–302 D. M. Smith and D. Main, Space Sciences Laboratory, University of

California at Berkeley; F. Marshall and J. Swank, Goddard Space Flight Center, NASA; W. Heindl, University of California at San Diego; and M. Leventhal, University of Maryland, College Park, report: “The RXTE Proportional Counter Array detected a new x-ray transient on Aug. 12.60 and 12.72 UT during slews across the galactic-center region. It was not seen on Aug. 3, 14, 20, or 28. The position is α = 17^{h}38^{m}59^{s}, δ = −30^{o}08′.4 (equinox 2000.0; 3′ error radius at 99-percent confidence). The 2–20-keV flux was 70 mCrab, and the spectrum could be fit with a hard (photon index −1.6) power law plus a 4-keV disk blackbody. The rms variability was 13 percent from 0.005 to 10 Hz. The power spectrum showed no flattening at low frequencies and no pulsations in this range. Although it is no longer bright in x-rays, prompt observations, particularly in the radio, are encouraged.”

COMET 78P/GEHRELS 2 Visual m_{1} estimates: Aug. 30.09 UT, 12.8 (J. Carvajal, Avila, Spain,

0.32-m reflector); Sept. 1.11, 11.9 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m refl.); 11.10, 12.4 (R. J. Bouma, Groningen, The Netherlands, 0.25m refl.); 12.06, 12.2 (M. Meyer, Frauenstein, Germany, 0.25-m reflector).

1997 September 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green